I decrease a certain number by two fifths of itself, and it becomes one more than half its original value. What is the number?
Answer by stanbon(75887)

I decrease a certain number by two fifths of itself, and it becomes one more than half its original value. What is the number?
----
Equation:
x - (2/5)x = (x/2)+1
-----
(3/5)x - (1/2)x = 1
----
(6+5)/10x = 1
-----
(11/10)x = 1
x = 10/11
--------------
Cheers,
Stan H.
------------